OpenStudy (alurahenderson):

What is the solution to the equation? 2q + 18 = –5q – 3 A. q = –5 B. q = –3 C. q = 3 D. q = 5

pooja195 (pooja195):

Isolate q

pooja195 (pooja195):

Start by combining the LIKE TERMS

OpenStudy (alurahenderson):

20q?

pooja195 (pooja195):

hmmm not quite We want to combine like terms this would mean subtracting and adding from the other side of the equation

pooja195 (pooja195):

Step 1: Add 5q to both sides. 2q+18+5q=−5q−3+5q

OpenStudy (alurahenderson):

Oh so -2 on both sides?

OpenStudy (alurahenderson):

I didn't see that

pooja195 (pooja195):

You can also do it that way :)

OpenStudy (alurahenderson):

22q and 8q?

pooja195 (pooja195):

\[\huge ~2q+18+5q=−5q−3+5q\]

pooja195 (pooja195):

whats 5+2=? whats -5+5?

OpenStudy (alurahenderson):

7 and 0

pooja195 (pooja195):

Good so now we have \[\huge~7q+18=-3\] any ideas on what the next step may be?

OpenStudy (alurahenderson):

-3 on both sides?

pooja195 (pooja195):

hmm no we are isolating q so subtract 18

pooja195 (pooja195):

From both sides

OpenStudy (alurahenderson):

I was gonna say -18 but I didn't think so

pooja195 (pooja195):

Well you were right :) -3-18=?

OpenStudy (alurahenderson):

0

OpenStudy (alurahenderson):

And -11

pooja195 (pooja195):

no try again.

OpenStudy (alurahenderson):

-21

pooja195 (pooja195):

good whats -21/7

OpenStudy (alurahenderson):

-3

pooja195 (pooja195):

Good so we get x=-3 :)

OpenStudy (alurahenderson):

Thank you again!!

pooja195 (pooja195):

:)